91,375.73J of heat is added to 1000 g of ice at -10 oC. How much ice melts?
Answer #1

Amount of heat required to 1000g of ice to increase the temp by 10 c =m*c*t=1*2030*10=20300 J

rest amount=91375.73-20300=71075.73 J

mL=71075.73

m=71075.73/334000=212.8 g

Amount of heat required to 1000g of ice to increase the temp by 10 c =m*c*t=1*2030*10=20300 J

rest amount=91375.73-20300=71075.73 J

mL=71075.73

m=71075.73/334000=212.8 g

Amount of heat required to 1000g of ice to increase the temp by 10 c =m*c*t=1*2030*10=20300 J

rest amount=91375.73-20300=71075.73 J

mL=71075.73

m=71075.73/334000=212.8 g